Find the quadratic function that fits the set of data points.
(-3,-5), (-2,-8), (6,-248)

Answer by stanbon(75887)

Find the quadratic function that fits the set of data points.
ax^2 + bc + c = y
---
Solve for a,b,c:
---
(-3,-5) gives: 9a -3b + c = -5
(-2,-8) gives: 4a -2b + c = -8
(6,-248)gives:36a+ 6b + c = -248
-------------------
I used a matrix function on my TI-84 to get:
a = -3
b = -18
c = -32
-----
Equation:
y = -3x^2 -18x - 32
=======================
Cheers,
Stan H.
=================